OpenStudy (anonymous):

8-5/8x>9 what is the solution set solve the linear inequality

OpenStudy (pfenn1):

Rearrange this inequality so x is on one side of the inequality and the numbers are on the other. Do you know how to do this?

OpenStudy (anonymous):

can you help me

OpenStudy (pfenn1):

\[8-\frac{5}{8x}>9\]Subtract 8 from both sides.

OpenStudy (anonymous):

is it 8-8

OpenStudy (pfenn1):

Close.\[\cancel8-\cancel8-\frac{5}{8x}>9-8 \]Which simplifies to \[\frac{-5}{8x}>1\]Do you know what to do now?

OpenStudy (anonymous):

no

OpenStudy (pfenn1):

Try to multiply each side of the inequality by 8x. What do you get?

OpenStudy (anonymous):

not sure

OpenStudy (pfenn1):

Did you try it?

OpenStudy (anonymous):

i do not know where to start

OpenStudy (pfenn1):

Let's try to multiply both sides of the inequality by x (instead of 8x). I am trying to get just x on one side of the inequality and all the numbers on the other side.\[x \times \frac{-5}{8x}>1 \times x\] What is the resulting inequality?

OpenStudy (anonymous):

1*x

OpenStudy (pfenn1):

Okay. That is what you get on the right side of the inequality. But what do you get on the left side of the inequality?

OpenStudy (anonymous):

not sure

OpenStudy (pfenn1):

what is \[x \times \frac{-5}{8x}\]

OpenStudy (anonymous):

the solution set

OpenStudy (pfenn1):

?

OpenStudy (anonymous):

x*-5/8x

OpenStudy (pfenn1):

\[\cancel x \times \frac{-5}{8\cancel x}>1 \times x\]The inequality now simplifies to \[\frac{-5}{8}>x\]

OpenStudy (anonymous):

so is that the solution-5/8>x

OpenStudy (pfenn1):

Yes. I think you need to work on how to rearrange equations and inequalities to find a solution.